### v3

# Survey of Natural Language Processing for Education: Taxonomy, Systematic Review, and Future Trends

Abstract
Natural Language Processing (NLP) aims to analyze text or speech via techniques in the computer science field. It serves the applications in domains of healthcare, commerce, education and so on. Particularly, NLP has been widely applied to the education domain and its applications have enormous potential to help teaching and learning. In this survey, we review recent advances in NLP with the focus on solving problems relevant to the education domain. In detail, we begin with introducing the related background and the real-world scenarios in education where NLP techniques could contribute. Then, we present a taxonomy of NLP in the education domain and highlight typical NLP applications including question answering, question construction, automated assessment, and error correction. Next, we illustrate the task definition, challenges, and corresponding cutting-edge techniques based on the above taxonomy. In particular, LLM-involved methods are included for discussion due to the wide usage of LLMs in diverse NLP applications. After that, we showcase some off-the-shelf demonstrations in this domain. At last, we conclude with six promising directions for future research, including more datasets in education domain, controllable usage of LLMs, intervention of difficulty-level control, interpretable educational NLP, methods with adaptive learning, and integrated systems for education. We organize all relevant datasets and papers in the open-available Github Link for better review https://github.com/LiXinyuan1015/NLP-for-Education.
Index Terms:
Natural language processing, educational NLP, educational applications, question answering, question construction, automated assessment, error correction, survey.
1 Introduction
FIGURE 1: We demonstrate the education scenarios in real-world and how NLP applications contribute to these scenarios.
Natural Language Processing (NLP) refers to the branch of computer science and artificial intelligence, that empowers computers the ability to understand text and spoken words as the way human beings can. It serves real-world scenarios. For example, NLP can help in healthcare applications. A wide range of NLP techniques are being utilized to anticipate the treatment process of the patients. Question answering [1, 2] and dialogue generation [3] help conduct medical diagnosis, administrative assistants for appointments and billing. Besides, NLP can help in commerce applications. There is a growing trend of incorporating NLP techniques into analytics and customer management platforms to help learn more about the relationship between the consumers and items via sentiment analysis [4].
The widespread collaboration of new technologies into the education domain revolutionize education in the future. In the education domain, NLP techniques can help education in the following ways:
  • NLP technologies provide automatic key information extraction which facilitates to highlight the key points. In the cases where teachers and students read passages, tutorials, and textbooks. The key points and knowledge should be annotated or noticed from unstructured text. Automatic key information extraction can assistant to identify the key notes which help them understand and comprehend during teaching and learning.
  • NLP technologies offer intelligent tutoring and recommendation. For example, when students are learning a concept, a system can help recommend relevant concepts, practices, and reading materials such that the students can find proper reading materials as references, which enables active learning and self-inspection for students.
  • NLP technologies help automatic problem solving. For example, in math or science courses, students may have extensive questions towards the exercises after the classes. To avoid repeat and overloading questions for teachers, NLP technologies can conduct reasoning and answer the questions automatically. Moreover, they can give explanations to the complex questions and intermediate solutions for the problems.
  • NLP technologies facilitate automatic smart content generation. For example, teachers have trouble in preparing the teaching materials from a rush of information on the Internet or textbooks. With the help of retrieval technologies, we can create a wide variety of digital learning materials, such as retrieving reading passages for students in a reading course. Similarly, given a passage in the textbook as the context and a concept as the correct answer, NLP techniques can quickly construct a multiple-choice question associated with the test question and the distractors. This could facilitate the teachers to make quick quizzes for classes.
  • NLP techniques enables personalized instruction. Students may come from different levels, learn different subjects and have difficulties in different concepts. In such a case, it is vital to track the learning trajectory and provide personalized instruction to the students. Various NLP models trained on different annotated data collected from different scenarios could contribute to the personalized education.
  • NLP technologies achieve AI-supported assessment and correction. For example, it is time-consuming for teachers to grade the essays and codes from students in linguistics or coding classes, by analyzing the textual contents of the assignments, NLP applications can figure out the errors in the assignment and return the corresponding correction to the students.
We display the education scenarios in real-world, relevant NLP techniques, and the potential contributions of them in Figure 1. These techniques are able to analyze students performance, allowing for more flexible learning, thus freeing up valuable time for teachers.
We notice there are a few surveys related to NLP for education. However, a line of surveys focus on a certain task. Specifically, Lan et al. [5] summarize the general advancement for knowledge base question answering tasks but the domain is not restricted to education, which maybe inapplicable for educational question answering. Bryant et al. [6] summarize the cutting-edge techniques of grammatical error correction, which is a typical tasks of educational NLP. Zhang et al. [7] carefully view the technical progress of math word problem solving tasks. Messer et al. [8] systematic review the task of code programming. However, the above surveys fail to identify the significant role and distinction of these tasks applied in education domain. Another line of surveys summarize in the shallow perspectives of concepts and definition [9, 10, 11], which are in a high-level and do not shed any light on the techniques and applications. Such that they cannot serve as the good references for researchers in NLP and computer science communities.
To bridge the gap, this survey provides a comprehensive overview of NLP for education from a technical point of view. (1) New taxonomy: we present a novel taxonomy for educational NLP and identify representative NLP tasks that can be applied for learning and comprehension as well as writing and assessment in education scenarios, and then discuss them in detail; (2) Comprehensive review: we collect recent papers published in well-recognized computer science venues and review their methods for solving the challenges of corresponding tasks. We further include LLM-based methods due to the wide usage in NLP applications; (3) Rich resources: we discuss the publicly-available datasets and demonstrations with their various attributes, which makes it easy for readers to get started for their researches.
For the rest of the manuscript, we first present the taxonomy of NLP in the education domain, which highlights four representative tasks and eight fine-grained sub-tasks in total. For each task, we review the datasets and techniques for the identified tasks, pointing out the major challenges and corresponding solutions. Then, we summarize the available demonstrations. At last, we conclude the survey with future directions.
FIGURE 2: An overview of NLP for education. The left section displays the systematic frame of relevant NLP techniques, including data source, processing and analysis techniques, and applications. The right section demonstrates the outcomes of diverse techniques.
2 Taxonomy of NLP in Education Domain
Systems built on the basis of NLP applications and equipped with NLP techniques could perform as empowered AI tutoring, which enhances the teaching and learning experience of students, teachers, parents, or tutors. With the systems, students are able to get access to private tutoring and extra support outside the classroom, which avoids increasing the workload of their teachers.
We first demonstrate an overview of NLP for education in Figure 2, where the systematic frame of relevant NLP techniques and their corresponding outcomes are displayed. The data sources are significant as the fundamental supports for subsequent analyses and applications. Generally, we could obtain educational data from textbooks, syllabus, exams, and curricula materials, which may contain behavioral data of students, exercise questions, and guidance signals of error correction and so on. From the perspective of teaching subjects, the data resources cover subjects like chemistry, math, linguistics, computer science and so on, which could support applications in different subjects. Given the raw data resources, we could conduct processing and analysis (e.g., knowledge graph construction [12], information extraction [13], information retrieval [14], knowledge alignment [15]) to extract the key information from massive data and organize them in structured knowledge graphs (e.g., EduKG [^1] https://edukg.cn/team), databases (e.g., WordNet [^2] https://wordnet.princeton.edu/documentation) and lists, which could be applied to various downstream applications. Upon the well-organized data, NLP tasks are applied (e.g., question answering, question construction, automated assessment, and error correction) to form high-level applications to serve teachers and students in real-world scenarios. Grammarly [^3] https://app.grammarly.com, Gauthmath [^4] https://www.gauthmath.com, CodeSubmit [^5] https://codesubmit.io, AutoQG  [^6] https://autoqg.net, and EduChat [^7] https://www.educhat.top/ are possible educational products and demonstrations developed based on these NLP applications.
This survey mainly focuses on the application-level. We highlight the most representative NLP applications with the taxonomy shown in Figure 3. Considering that NLP techniques play significant roles in different phases of education, on the one hand, NLP could serve as a teaching assistant for learning and comprehension. When students learn the curriculum, the following NLP tasks could simplify the difficulties of learning.
  • Question Answering (QA), as a typical application, could not only provide tutorials to the homework for students but also save time for overloaded educators by providing feedback on questions. Based on the different goals of question answering, textbook question answering and math word problem solving are presented as the two typical subcategories of question answering. These two tasks both aim to generate answers from the given contexts automatically. They can be widespread in the education of science subjects.
  • Question Construction (QC) is to generate questions as well as the distractors given the correct answers, which helps the teachers create a wide range of test questions using pre-existing learning resources. Generally, a good test question may consist of a question and multiple candidate answers. Therefore, question generation and distractor construction are two key sub-tasks to be reviewed.
On the other hand, NLP improves writing and assessment for students. When given constructive criticism and writing prompts, students are likely to go deep into the topic. However, teachers sometimes have difficulties responding thoughtfully to students due to time limitation. But they will benefit from the following tasks.
FIGURE 3: The taxonomy of NLP for Education. The NLP applications associated with the well-defined tasks and applicable subjects are displayed.
  • Automated Assessment (AA) aims to grade assignments for students automatically. Language grading is a representative scenario of AA, which includes automated essay scoring for natural language learning and automated code scoring for program language learning. For essay scoring, grammar, spelling, and sentence structure are key aspects that need to be considered. For code scoring, the correctness of the code, style aspects, code quality are metrics to be viewed.
  • Error Correction (EC) for language are useful to improve language learning via explicating the correct edits. Still, taking language learning as an example, grammatical error correction and code error correction are canonical sub-tasks for EC. The methods are not simply required to provide error-free outputs based on the ill-posed inputs, more detailed instructions should be provided to shed light on the reasons or knowledge points to the errors.
Please be aware that there are some other NLP tasks that are related to education. For example, tasks like text rewriting [16] and controllable text generation [17] can also help teachers to prepare suitable teaching materials. In this survey, we mainly focus on the highlighted tasks, which are widely explored NLP tasks in education domain. Our discussion is carried out based on the above taxonomy.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=left, calign=edge midpoint, , [NLP Applications for Education, text width=2.7cm, for tree=fill=white!20 [Textbook Question Answering, text width=1.8cm, for tree=fill=white!20 [Comprehend the multi-modal contexts and complex questions, text width=2.8cm, for tree=fill=white!20 [ Propose deep learning-based methods combining RNN and CNN [18, 19]; spatial attention [20, 21, 22, 23]; compositional strategies [24]; bilinear pooling schemes [25, 26]., text width=6.5cm, node options=align=left ] ] [Supplement domain-specific knowledge, text width=2.8cm, for tree=fill=white!20 [ Pre-train models with domain knowledge [27, 28]; diagrams parsing enhancement [29, 30, 31]., text width=6.5cm, node options=align=left ] ] [Strengthen zero-shot capabilities, text width=2.8cm, for tree=fill=white!20 [ Collaborate LLMs with textbooks [1, 32, 33, 34]; develop LLMs-based multi-agents [35, 2]., text width=6.5cm, node options=align=left ] ] ] [Math Word Problem Solving, text width=1.8cm, for tree=fill=white!20 [Understand the gap between natural language and abstract expression, text width=2.8cm, for tree=fill=white!20 [Propose seq2seq [36, 37, 38, 39] framework enriched with structured encoders and decoders [40, 41]; reinforcement learning [37, 42]; knowledge distillation [43]., text width=6.5cm, node options=align=left ] ] [Solve MWPs under zero/few-shot setting, text width=2.8cm, for tree=fill=white!20 [ Propose LLMs-enhanced MWP solvers [44, 45, 46, 47, 35]; Develop LLMs-based multi-agents [48]., text width=6.5cm, node options=align=left ] ] ] [Question Generation, text width=1.8cm, for tree=fill=white!20 [Identify the key statement from the context and generate the corresponding question, text width=2.8cm, for tree=fill=white!20 [ Conduct rule-matching [49]; develop deep learning-based methods with attention [50]; lingtuistic features [51, 52, 53]; multi-modal fusion [54]; multi-task learning [55]; reinforcement learning [56]., text width=6.5cm, node options=align=left ] ] [Generate more flexible questions for adaptive education, text width=2.8cm, for tree=fill=white!20 [ Propose difficulty-control QG methods [57, 58, 59]; syllabus-aligned methods [60, 61]; knowledge-tracking QG methods [62, 63]., text width=6.5cm, node options=align=left ] ] ] [Distractor Generation, text width=1.8cm, for tree=fill=white!20 [Design context-consistent and deceptive false answers, text width=2.8cm, for tree=fill=white!20 [ Propose rule-based methods [64]; feature-based methods [65]; generation-based methods [66, 67]; ranking-based methods [68, 69, 70]., text width=6.5cm, node options=align=left ] ] ] [Automated Essay Scoring, text width=1.8cm, for tree=fill=white!20 [Model different facets of the essay, text width=2.8cm, for tree=fill=white!20 [ Design various features (e.g. length-based features [71]; lexical features [72, 71]; prompt-relevant features [73]; readability features [71]; syntactic features [71, 74]) and solve via regression and ranking [75] with enhanced training procedure [76]., text width=6.5cm, node options=align=left ] ] [Sparse annotation of essays, text width=2.8cm, for tree=fill=white!20 [ Decouple the domain-dependent and domain-independent features and aggregate heuristic supervision signals [77, 78, 79]; design enriched annotations [80, 81]., text width=6.5cm, node options=align=left ] ] ] [Automated Code Scoring, text width=1.8cm, for tree=fill=white!20 [Understanding of the long-dependency of code, text width=2.8cm, for tree=fill=white!20 [ Conduct dynamic analysis to assess the execution of the code [82, 83, 84, 85]; static analysis to predict scores with the designed features [86, 87, 88, 89]; hybrid methods combine both [90]., text width=6.5cm, node options=align=left ] ] ] [Language Error Correction, text width=1.8cm, for tree=fill=white!20 [Identify and correct grammar errors accurately, text width=2.8cm, for tree=fill=white!20 [Design manual rules and classification-based models [91, 92]; neural network models like seq2seq models [93, 94, 95, 96, 97] and seq2edit models [98, 99, 100, 101]., text width=6.5cm, node options=align=left ] ] ] [Code Error Correction, text width=1.8cm, for tree=fill=white!20 [Identify and correct errors in the long-distance dependency of the code, text width=2.8cm, for tree=fill=white!20 [Develop language modeling [102, 103]; methods with abstract syntax tree [104, 105]; denoising strategies [106]; use identifier tagging and comment information [107]; grammar constraints [108]., text width=6.5cm, node options=align=left ] ] ] ]
FIGURE 4: The detailed techniques of NLP applications. The hierarchical structure is arranged with: NLP applications $\rightarrow$ Challenges $\rightarrow$ Solutions.
3 Review of Techniques
To help readers catch up with cutting-edge research on these highlighted tasks, in this section, we provide a comprehensive review of the evaluated resources and techniques for these tasks, showing emphasis on the development of techniques and solutions to the key challenges. The review is arranged in a tree in Figure 4 for a quick lookup.
3.1 Question Answering
FIGURE 5: The examples of Question Answering in education domain. The red box denotes the given context and the green box denotes the required prediction for the tasks. The displayed examples are extracted from SCIENCEQA and GSM8K, respectively.
Even though QA has been explored as a popular NLP task for decades. Educational QA requires a system to answer the domain-specific questions, related to natural science, social science, language science and so on [109]. These systems act as assistants to help answer difficult questions such that the students could learn from them and boost their performance. Question answering in different subjects may be presented in different formats. In Figure 5, we display two typical examples of question answering in the education domain. To answer the displayed questions, the systems need domain-specific knowledge like mathematics, and science. Furthermore, the contexts of the questions are presented as text and images. Moreover, the output of the questions can presented as an entity or a math expression. In this section, we will present cutting-edge techniques for Textbook QA and MWP solving.
3.1.1 Textbook Question Answering
Textbook Question Answering (TQA) is a task that requires a system to comprehensively understand the multi-modal information from the textbook curriculum, spreading across text documents, images, and diagrams. The major challenge of textbook question answering is to comprehend the multi-modal domain-specific contexts as well as the questions, and then identify the key information to the questions.
Datasets. Kembhavi et al. [110] introduced TQA dataset, which aims to evaluate a system that contains multi-modal contexts and diverse topics in the scientific domain. Similar datasets such as AI2D [29], DVQA [31], and VLQA [111] have been constructed to support researches on multi-modal reasoning in the scientific domain. However, these datasets lack of annotated explanations for the answers in the form of supporting facts. SCIENCEQA [1] is a large-scale textbook question answering dataset with annotated lectures and explanations. The dataset is collected from elementary and high school science curricula, which contain diverse science topics including natural science, social science, and language science. A recently released TheoremQA dataset further covers textbook questions at the university level [2]. Besides the scientific domain, there are some datasets focusing on the medical domain. MEDQA [112] and MedMCQA [113] are two medical QA datasets encompassing a broad spectrum of healthcare topics deriving from both real-world scenarios and simulated exams.
Methods. From the technical perspective, textbook question answering is visual question answering in nature [114, 115, 116]. Conventional VQA studies employ RNN to encode the question and CNN to encode the image [18, 19], respectively. Then the multi-modal information is fused together to comprehend the questions. Besides, other methods employing spatial attention [20, 21, 22, 23], compositional strategies [24], and bilinear pooling schemes [25, 26] are proposed to boost the performance of VQA.
While there are significant similarities between VQA and textbook question answering, textbook question answering differs in the domain-specific knowledge required for the accompanying context and the innovative inclusion of diagrams and tables. To bridge the gap, Ram et al. [27] proposed a pre-training schema tailored for question answering. Specifically, by masking recurring span selection and selecting the correct span in the passage, their method boosts the performance for textbook question answering when only a hundred examples are available in the specific domains. Adversarial training framework is also adapted to domain generalization [28], which facilitates the question answering models to learn domain-invariant features. Xu et al. [117] introduced a novel Pre-trained Machine Reader as a retrofit of pre-trained Masked Language Models (MLMs), which resolves the discrepancy between model pre-training and down-stream fine-tuning of specific domain MLMs. To understand the diagrams and tables, graph-based parsing methods are developed to extract the concepts from diagrams [29] by converting a diagram to a diagram parse graph. Optical Character Recognition (OCR) is employed to identify the chart-specific answers from the charts, which is further aligned with the questions [30, 31].
Recently, Large Language Models (LLMs) have shown impressive capabilities in various NLP tasks, researchers start applying LLMs to textbook question answering. Lu et al. [1] first applied GPT-3 to solve the textbook question answering. To apply the full advantage of LLMs to image modality, they use caption models to translate visual information into the language modality and prompt LLMs to perform Chain-of-Thought. This inspires the follow-up study [32], distilling the training signals from the LLMs to a smaller model. To inject general knowledge from the LLMs to the specific domains, Wang et al. [33] augmented queries for LLMs on textbook question answering by retrieving evidences from textbooks, such that the response of the LLMs would likely to be authorized by the textbooks. Liu et al. [34] extended the instruction tuning to a multi-modal field, which results in a multi-modal model that connects a vision encoder and an LLM for general-purpose visual and language understanding. More recent studies [35] develop compositional reasoning with LLMs as a planner that assembles a sequence of tools (e.g., Hugging Face as image captioner, Github as text detector, Web Search as bing search, and Python as program verifier.) to execute and answer the textbook questions. Instead of solving the questions with given contexts, Chen et al. [2] introduced a system, that first retrieves the theories from the open domain and then answers the question based on the retrieved theories.
Dataset Science subject Source level Context #Q
TQA life/earth/physical grade 6-8 image/ 26262626K
science science curricula diagram
Geometry3K geometry grade 6-12 digital libraries image/diagram 3333K
AI2D science grade 1-6 image/diagram 5555K
SCIENCEQA natural/social/ grade 1-12 image/diagram 21212121K
language science science curriculum text
MedQA medicine professional medical text 40404040K
board exam
MedMCQA medicine simulated exams text 200200200200K
TheoremQA math/physics/EE/CS university-level exam text $800800$
Dolphin-18K math grade school text 18181818K
DRAW-1K math algebra text 1111K
Math23K math grade school math text 23232323K
MathQA math grade school math text 37373737K
ASDiv math grade school math text 2222K
GSM8K math grade school math text 8888K
IconQA math K-12 math text/image 107107107107K
TABMWP math grade school math text/table 38383838K
TABLE I: Comparison of textbook question answering and MWP solving datasets. The first section includes datasets of Textbook QA. The second section includes datasets of MWP solving. “#Q” denotes the number of questions.
3.1.2 Math Word Problem solving
Math Word Problem (MWP) solving is a specific type of question answering, aiming to convert a narrative description to an abstract expression. This task is challenging because there remains a wide semantic gap to parse human-readable words into machine-understandable logic.
Datasets. Dolphin-18181818[118] is an early large-scale MWP solving dataset, which contains over 18,0001800018,00018 , 000 annotated math word problems of elementary mathematics. DRAW-1111[119] is introduced as a testbed to algebra word problem solvers. Math23232323[36] is a Chinese MWP solving dataset. Subsequently, MathQA [120] and ASDiv [121] further extend the scale and diversity of the existing datasets, which makes it particularly suited for usage in training deep learning models to solve MWP problems. Cobbe et al. [122] introduced a more challenging dataset GSM8888K, which is complicated in high linguistic diversity. IconQA [123] extends MWP solving to abstract diagram understanding and comprehensive cognitive reasoning by involving diagrams in problems. Moreover, TABMWP [124] includes tables to enhance the difficulty of the MWPs, which requires the table parsing and understanding capabilities [124]. As we can see, there is a set of MWP datasets that examine various mathematics capabilities. With the development of this task, the evaluated datasets become increasingly complicated in linguistic diversity and multi-modalities.
Methods. The investigation of MWP solving can be traced back to 1960196019601960[125]. Early work leverages manually crafted rules and schemas for pattern matching, where human intervention is heavily relied on. Subsequently, various strategies of feature engineering and statistical learning are proposed to boost the performance. With the development of techniques and the emergence of large-scale datasets, deep learning-based methods have been widely applied to MWP solving. The majority of the methods follow an encoder-decoder framework [36], which first encodes the narratives into dense vector space, then generates the mathematical expression token as the output sequence. To improve the generation of complex solution expressions, some methods are proposed to formulate a mathematical expression as an abstract tree [37, 38, 39, 43]. To improve the understanding to the narratives, some studies focus on improving the encoding of the narratives by re-organizing the textual contexts with structure [40, 41]. Due to the weak supervision caused by the lack of annotated expressions, multiple studies [37, 42] employ reinforcement learning to solve MWPs. Also, some researchers [126] apply knowledge distillation to MWP solving tasks, which learns a smaller model from a large pre-trained generic model.
As the MWP solving task is the desirable testbed for evaluating the reasoning capabilities, a series of studies have been developed to discuss the abilities of LLMs for solving MWPs. Researchers find that prompting the LLMs with effective planning strategies could activate the potential of LLMs, which results in impressive results on MWP solving under the zero/few-shot scenarios. Chain-of-thought is a typical promoting strategy that is meant to decompose a complicated MWP to sub-questions such that LLMs could conduct reasoning step by step [127]. Follow-up methods improve it by designing more interactive prompting or constructing more efficient demonstrations [44, 45, 46, 47, 35]. With the emergence of the idea of leveraging LLMs as an agent [128], researchers [48] delegate the computation to a program interpreter, which can be conducted by the LLMs and executed by a Python interpreter. In this case, the reasoning step can be verified in an iteration manner.
3.2 Question Construction
Question Construction (QC) aims to automatically construct questions from a given context, which performs a significant role in education [129]. Usually, Multiple Choice Questions (MCQ) are common question types for a test quiz. The construction of MCQ consists of two parts: Question Generation (QG) and Distractor Generation (DG). The former generates questions based on a given context, while the latter generates distractors to complement the correct ones. In Figure 6, we display a typical example of QG and DG in the education domain. In this section, we introduce the relevant datasets and methods for QG and DG in detail.
FIGURE 6: The examples of Question Construction. The displayed examples are extracted from SciQ.
3.2.1 Question Generation
The goal of QG is to automatically generate a question from a given sentence or paragraph, with the assistance of answer information when available. The challenge lies in identifying the key statement based on the context with or without the utilization of the answer (namely answer-aware QG or answer-agnostic QG [130]) and generating a question based on the statement.
Datasets. QG and QA are dual tasks [131]. They both require reasoning between questions and answers. As a result, some MCQ datasets initially designed for QA tasks, including SciQ [132], RACE [133], and FairytaleQA [134], are also leveraged for researches on QG [135, 136, 137, 138]. Besides, there are some datasets particularly constructed for QG, such as LearningQ [139], KHANQ [140], and EduQG [141]. These datasets extend across various subjects (e.g., science, medicine) and various education levels (e.g., preliminary school, middle school, and university.), the details of which are displayed in Table II. Regarding these datasets, LearningQ is collected from instructor-crafted learning questions, which contain cognitively demanding documents for QG that requires reasoning, but the absence of answers limits its application to answer-aware QG. KHANQ provides (context, prompt, question) triples collected from questions about the online courses raised by students. The prompt, which indicates the background knowledge of students or presupposition of the questions, provides hints on the information that should be brought into focus during question generation. EduQG is a high-quality MCQ dataset generated by educational experts and questions are linked to their cognitive complexity, which makes this dataset move closer to reality.
Dataset Science subject Source Question type Answer type #Q
SciQ science online science normal MC& 13.713.713.713.7K
textbooks extractive
RACE linguistic secondary school normal& MC& 100100100100K
English exams cloze extractive
FairytaleQA literature grade K-8 normal extractive 10101010K
narrative stories
LearningQ computing/ online education normal - 231231231231K
science/business/ platform
humanities/math
KHANQ science online education normal extractive 1111K
platform
EduQG politics/physiology/ online textbooks normal& MC& 3333K
biology/business/ cloze extractive
law/sociology/
psychology/history
MCQL biology/physics/ website normal MC 7.17.17.17.1K
chemistry
Televic math/health/ online education normal MC 62626262K
history/geography/ platform
science
TABLE II: Comparison of QG and DG datasets. The format of the question is categorized as either normal (e.g., interrogative sentence.) or cloze (e.g., fill-in-the-blank.), and the answer is categorized as either multiple-choice (MC) or extractive (parts of the text are considered as the answer) [141].
Methods. Early methods for general QG primarily relied on rule matching [49]. Du et al. [50] first applied the Seq2Seq model with attention to automatic QG. Built upon this framework, methods such as integrating linguistic features [51, 52, 53], training a multi-modal QG model [54], leveraging multi-task learning [55] and reinforcement learning [56] have also been augmented to optimize the QG models. With the advent of language models, the BERT model and LLM GPT-3 have been applied to QG [142, 143]. Besides, regarding answer information in the QG system, researchers proposed to encode the answer together with the context [144], utilize the position information of the answer in the context [145, 146], and achieve answer-aware QG, while using summaries of the texts as input can serve as a strategy for answer-agnostic QG as it can help the modeling of topic salience, which denotes the importance of specific words or phrases in context [130].
Importantly, there are some unique notes when we adopt QG to the education domain. First, controlling the difficulty levels of the generated questions is vital for improving the efficiency of education. Gao et al. [57] proposed a method for difficulty-controllable QG, and the difficulty of a question is judged according to whether a well-trained QA model is able to answer it correctly. Then Cheng et al. [58] redefined the difficulty level as the number of inference steps required for a question to answer. Recently, Uto et al. [59] considered the difficulty associated with learners’ ability, which is measured by the item response theory [147]. Second, it is important to align generated questions to the syllabus, which benefits to identify the focus of the test quiz. Hence, multiple studies [60, 61] train binary classifiers or ranking models to judge whether the generated question is related to certain entities and concepts or pedagogical contents with the objective of maximizing both relevance and diversity. Third, customized education is in high demand and we notice there are a few studies focusing on generating personalized questions for students. Srivastava et al. [62] developed a knowledge tracking model that uses students’ answer histories to predict their performance on new questions, fine-tuning it with an autoregressive language model. Subsequently, Wang et al. [63] introduced the few-shot knowledge tracking model to predict question difficulty by incorporating sequences of student states and questions.
3.2.2 Distractor Generation
As we mentioned above, DG is another component for question construction, which aims to generate distractors with the given context, question, and correct answer. A good distractor should be relevant to the context, grammatically coherent to the interrogative of the question, and deceptive as a false answer [132].
Datasets. The aforementioned MCQs for QG tasks such as SciQ and RACE can also be leveraged for DG tasks. Specifically, the distractors in SciQ are designed by crowd workers with reference [132]. Instead of providing entities and concepts as distractors, the distractors of RACE are at sentence-level [148], which are collected from the secondary school English exams. Alongside the datasets mentioned before, there are other MCQ datasets especially designed for DG. The data of MCQL [68] is crawled from the web and provides distractors that are either phrases or sentences, covering multi-subjects. Besides, Televic [149] is collected through Televic Education’s platform with distractors manually crafted by experts, featuring distractors across multiple languages and domains.
Methods. Early studies for DG are rule-based and feature-based, which rely on linguistic rules [64] and specific features like word frequency [65] or diverse similarity measurement between distractors and correct answers [68]. With the rapid development of deep learning, the construction of distractors can be divided into two lines of mainstream: generation-based and ranking-based categories [70]. Generation-based methods automatically generate distractors token by token and ranking-based methods regard DG as a ranking task for a pre-defined distractor candidate set. Amongst the generation-based methods, Gao et al. [66] proposed a hierarchical model with dual attention for sentence and word relevance and irrelevant information filtering to generate distractors that are semantically consistent with and traceable in the context. Besides, Bitew et al. [67] employed LLMs for DG. They feed relevant questions as in-context examples to generate more plausible distractors. As for the ranking-based method, Liang et al. [68] proposed a two-stage framework: the first stage trains a model as a filter to reduce the candidate size, and the second stage trains another model for ranking the candidate. However, the two-stage framework encounters error accumulation, and some studies have discarded the first stage to address this issue, but led to reduced efficiency [69]. As follow-up studies, Wang et al. [70] proposed an end-to-end model for retrieval in the distractor candidates with a balance of efficiency and high performance.
3.3 Automated Assessment
FIGURE 7: The examples of AES and ACS tasks. The displayed examples are extracted from ASAP dataset and paper [150], respectively.
Automated assessment is a widely-investigated task in the education domain, which is helpful in reducing the burden of teachers in various writing tasks. Regarding different objects to be scored, different criterion is considered. For instance, in Figure 7, scoring an essay in a language curriculum requires the measurement of particular essay quality such as persuasiveness and organization. For a piece of code in the program curriculum, it is more important to measure the correctness of the code, the completeness of the comments, and quality issues. Therefore, we discuss the techniques for Automated Essay Scoring (AES) and Automated Code Scoring (ACS) in the following subsections.
3.3.1 Automated Essay Scoring
AES aims to develop a system to score an essay automatically, where the written text is given as the input and the system summarizes its quality as a score. The challenge of this task lies in understanding the content of the essay by modeling the different facets of the text. Until now, a series of methods have been developed to automatically score the essays.
Dataset Essay Types Source Language Exp #Q
CLC-FCE A/N/C/S/L non-native: ESOL test English holistic 1111K
ASAP A/R/N native: grades 7-10 exam English holistic 17171717K
TOEFL 11 A non-native: TOEFL test English holistic 1111K
ICLE A non-native: English organization/clarity/ 4444K
undergraduate exam adherence/persuasiveness
HSK A/N non-native: HSK test Chinese holistic 10101010K
TABLE III: Comparison of AES datasets. The essay types contain: argumentative (A), response (R), narrative (N), comment (C), suggestion (S) and letter (L). “Exp” denotes the type of scores annotated. “#Q” denotes the number of questions.
Datasets. The Cambridge Learner Corpus-First Certificate in English exam (CLC-FCE) [151] is a widely used dataset for AES tasks. It provides both the holistic score and the manually tagged linguistic error types such that it makes possible to evaluate the instructional feedback of the developed AES systems. It is worth noting the dataset is annotated with the prompt of the essay, which indicates the topic of the essay. The Automated Student Assessment Prize (ASAP) is a Kaggle competition released for AES. A similar dataset TOEFL 11 corpus [152] contains essays from TOEFL exams, and it comes with a three-level proficiency annotation, which facilitates the AES research with the consideration of the difficulty level. Instead of holistic scores, fine-grained instructional annotation is needed for dimension-specific essay scoring research. International Corpus of Learner English (ICLE) and Argument Annotated Essays (AAE) [153] are two datasets containing essays annotated with the perspectives of structure organization, thesis clarity, prompt adherence, and argument persuasiveness. There are a few AES datasets, which are designed for other languages such as Chinese [154]. We display the details of the above datasets in Table III.
Methods. Due to the availability of the annotation and wide applicability of the commercial value, the majority of the studies usually predict a single score for an essay. The early AES systems can be formulated as a regression task, a classification task, or a ranking task. Specifically, a series of features such as length-based features [71], lexical features [72, 71], prompt-relevant features [73], readability features [71], syntactic features [71, 74] and so on, are first extracted from the essay, then off-the-shelf learning algorithms are leveraged to fit the annotated scores [155]. For regression, linear regression, support vector regression, and sequential minimal optimization are employed. For classification, SMO, logistic regression, and Bayesian networks are utilized. For ranking, SVM ranking can be applied. Subsequently, researchers propose neural networks with different architectures, including CNN, LSTM, and attention mechanism, to encode the essay and eventually predict the scores. Particularly, the document structure can be captured via hierarchically encoding the sentences from the word level to the sentence level. The supervision signals can be formulated with the combination of regression and ranking [75]. Hybrid methods are proposed to integrate the handcrafted features in deep learning-based methods [156]. Zeng et al. citezeng:aaai2023 applied curriculum learning to the AES training procedure.
Besides, there are some other challenges existing in the AES tasks. Due to the sparse annotation of the essays in the target domains, multiple studies [157, 158, 159] design neural networks decoupling the domain-dependent and domain-independent features, to facilitate the models to learn knowledge from different domains. Other methods aggregate multiple heuristic signals as supervision [79], enhance the features with capsule networks [78], and conduct interpretable essay scoring systems [77]. Regarding the evaluation of the essays, multiple dimensions are considered in a few studies such as cohesion, syntax, vocabulary, phraseology, grammar, and conventions[80, 81], which is able to construct more diverse evaluation metrics.
3.3.2 Automated Code Scoring
Similar to AES, ACS grades the score of a code snippet on various dimensions. Comparing with essay grading, code grading is more complicated in understanding the long-dependency inside a code snippet and the relationship between identifiers.
Datasets. We notice that there are few publicly-available ACS datasets due to the privacy issue. The source data contain submissions in programming courses that can be collected on the academy platforms such as JetBrains Academy, MOOC, tutoring sessions and so on [85, 160, 150].
Methods.
Code correctness is a vital factor in judging the quality of a code snippet, which can be quickly evaluated using compilers and interpreters. Therefore, dynamic analyses evaluate a code snippet by its performance of execution. By executing on a set of test cases, the execution results are supposed to be the same as the excepted results. In this way, many dynamic analysis tools for detecting code quality issues are developed, such as AutoStyle [82], WebTA [83], Refactor Tutor [84] and Hyperstyle [85]. On the contrary, some researchers apply static analysis approaches to assess a code snippet, where static features are extracted from the code without executing it. The static features include complexity, comments, number of lines, number of operators, number of operands and so on. The static analysis approaches leverage these features to identify the quality of code [86]. Machine learning algorithms also have been applied to ACS. Clustering, breadth-first search, and neural network approaches are utilized to track the logic errors in code [87, 88, 89]. Besides, Queiros et al. [90] proposed hybrid approaches combining static and dynamic approaches.
Instead of directly scoring a code snippet based on the extracted features, there are a series of studies focusing on evaluating the contextual embedding of a code snippet, which benefits to measure the distance between the evaluated code and the correct code for ACS [161]. The embedding of variable and method identifiers utilizes local and global context[162], abstract syntax trees [163], and AST paths [164, 165]. Later, Feng et al. [166] and Kanade et al. [167] proposed CodeBERT and CuBERT, respectively. They model code beyond the simple tokenization.
3.4 Error Correction
Compared with AA, EC requires systems to provide explicit edits and corresponding revisions to the errors. Similarly, language learning, as an important educational scenario, contains applications of Grammatical Error Correction (GEC) and Code Error Correction (CEC) for natural language learning and machine language learning. In terms of the techniques of GEC and CEC tasks, the CEC differs from GEC due to their different syntax. And we display two typical examples of them in Figure 8.
FIGURE 8: The examples of GEC and CEC tasks. The examples are extracted from BEA-2019 and Bugs2Fix, respectively.
3.4.1 Grammatical Error Correction
GEC tasks are designed to identify and correct grammatical errors in text, including spelling errors caused by phonological confusion and visual confusion, as well as grammatical errors caused by incorrect use of grammatical rules. For the GEC task, how to correctly identify and correct grammar errors is difficult, and how to avoid over-correction is challenging.
Datasets. Lang-8 is one of the most widely used datasets for language error correction, covering 80 languages, but its distribution of languages is very skewed, with Japanese and English being its most prevalent languages [168]. CLANG-8 [168] is a clean dataset based on Lang-8, fixing some mismatches between data sources and targets in Lang-8. CoNLL-2014 [169] and BEA-2019 [170] are both commonly used English language error correction datasets that refine different types of syntax errors. In addition, with the popularity of GEC tasks, more and more non-English datasets have been proposed and widely used by researchers. SIGHAN [171] is an early proposed data set for Chinese spelling check, which only contains grammatical errors such as pronunciation and shape errors of characters. The CTC [172] dataset not only contains spelling errors, but also introduces grammatical error and semantic error, which further improves the error types of the Chinese grammatical error dataset. The FCGEC [173] and FlaCGEC [174] datasets have added more linguistic annotations, introduced more granular grammatical rules, and covered 28 and 210 error types, respectively, presenting new challenges to the task of language error correction. At the same time, there are many GEC datasets for low-resource languages. There are dataset GECCC [175] developed in Czech, dataset RULEC-GEC [176] developed in Russian, dataset Falko-Merlin [177] developed in German, dataset COWS-L2H [178] developed in Spanish, dataset UA-GEC [179] developed in Ukrainian and dataset RONACC [180] developed in Romanian.
Methods. Early work to handle language error correction tasks is mostly achieved by manually defining rules or building classification-based models, with modest performance [91, 92]. With the widespread use of neural networks, Seq2seq [181] model has become the mainstream method to solve the task. It treats grammatical error correction as a monolingual machine translation task, in which source sentences containing grammatical errors are tagged one by one to be translated into the correct target sentences. The Seq2seq methods are mainly built based on RNN, CNN, or Transformer framework. BART [93], T5 [94], EBGEC [95], SynGEC [96] and CSynGEC [97] are good models in practice. After that, Sequence-to-editing (Seq2edit)  [182] methods have gradually emerged. It treats the GEC task as a sequential labeling problem, marking the text spans with appropriate error tags, leaving the rest of the text unchanged, and generating one edit operation per prediction. The commonly used efficient Seq2edit models include PIE [98], EditNTS [99], LaserTagger [100], GECToR [101] and so on. Recently, the TemplateGEC [183] model proposed by Li et al. integrates the two frameworks of Seq2seq and Seq2edit, making use of their ability in error detection and correction, alleviating the problem of over-correction of seq2seq model to a certain extent, enhancing the effectiveness and robustness of the model. As a result, it gains impressive results in GEC tasks. Even though LLMs are able to solve GEC tasks in nature, researchers found that they encounter severe issues of over-correction [184].
Dataset Language Source Exp #Q
LANG-8 Multi LANG-8 Website 1M
(80 languages)
CLANG-8 En, GE, RU LANG-8 2.6M
CoNLL-2014 En NUCLE [185] 58K
BEA-2019 En NUCLE, FCE, W&I, 686K
LOCNESS, LANG-8
SIGHAN Zh - 12K
CTC Zh Texts written by native speakers 218K
FCGEC Zh Chinese examinations 41K
FlaCGEC Zh HSK [186] 13K
GECCC CS Essays, website texts 83K
RULEC-GEC RU - 12K
Falko-MERLIN GE MERLIN, FalkoEssayL2, 24K
FalkoEssayWhig
COWS-L2H ES Essays 12K
UA-GEC UK - 20K
RONACC RO TV and radio shows 10K
Defects4J Java Open source programs - 357
ManyBugs C Open source programs - 185
IntroClass C Student programs - 998
QuixBugs Java, Python Quixey Challenge  [^8] https://gigaom.com/2012/01/19/quixey-challenge/ - 40
Bugs2Fix Java Github - 2.3M
CodeReview Multi Github - 642K
(9 languages)
CodeReview-New Multi CodeReview, code reviews - 15K
(16 languages)
TABLE IV: Comparison of GEC and CEC datasets. “Exp” denotes whether the annotation contains the explicit error types. “#Q” denotes the number of sentences in the dataset.
3.4.2 Code Error Correction
Code Error Correction (CEC) fixes a buggy code snippet to make it error-free and more coherent. It is a challenging task as it requires the model to have a sufficient understanding of the code in long-distance dependency and make proper corrections to the erroneous code spans.
Datasets. Defects4J [187], ManyBugs [188] and IntroClass [188] are some of the earlier datasets proposed for CEC tasks, focusing on Java and C languages respectively, but with small amounts of data. Lin et al. [189] proposed the first multi-lingual parallel corpus of CEC benchmarks, namely QuixBugs[189], which contains $4040$ buggy code snippets translated into both Java and Python. Xia et al. [190] combined the aforementioned Defects4J, ManyBugs, and QuixBugs datasets into a single CEC evaluation dataset to test the effect of PLM on the CEC task. Nowadays, the widespread use of collaborative software development platforms such as Github [^9] https://github.com/ makes it possible to collect and analyze code review data from open source projects. GitHub’s Forge provides a large amount of history-changing and bug-fixing commits from software projects, which can be the resource of CEC datasets in nature. Bugs2Fix [191] is one such dataset mined from GitHub bug-fixing commits, containing 2.3M bug-fix pairs (BFPS). CodeReview [192] is also a large-scale dataset of code changes and code reviews collected from open source projects, involving nine of the most popular programming languages. Because CodeReview has some low-quality code review data, and the data distribution in its test set is very similar to the data distribution in the pre-training and fine-tuning datasets, there is a potential risk when training with a model such as ChatGPT, which can affect the evaluation of the model’s generalization ability. Therefore, Guo et al. [193] create a new code review dataset CodeReview-New [193] based on it. By designing strict filtering rules to screen out low-quality data in CodeReview, they ended up retaining data from 467 repositories out of an initial 2,029.
Methods. The CEC and GEC tasks share similar principles and we can formulate both of them as sequence modeling, neural machine translation, and language modeling [102, 103]. Nevertheless, there are still some fundamental differences in their technical implementations. GEC tasks are typically solved by representing text as word embeddings or other forms of semantic vectors. While CEC tasks often require models to have a better understanding of long contexts, existing techniques capture the structural and semantic information of code by representing it as a sequence of symbols or an Abstract Syntax Tree (AST) [194]. Therefore, several pre-trained language models are trained using a large corpus of programming languages. CuBERT [167] is a pre-trained model for CEC tasks, which models only the Python language and is mainly applied to discriminative tasks of code understanding. CodeBERT [195] proposed in the same period is also modified on the basis of the BERT model, which is a multilingual programming language model that can handle dual-modal data (i.e. programming language (PL) and natural language (NL)). It supports modeling in 6 programming languages (i.e. Python, Java, JavaScript, PHP, Ruby, and Go) and combines pre-training tasks for replacement tag detection with hybrid objective functions, enabling the model to handle training using NL-PL data pairs. Both CuBERT and CodeBERT ignore the structural information of code. In response to this issue, Guo et al. [104] proposed the GraphCodeBert model to represent source code information based on AST data flow, so as to reflect deeper code information. As AST has a tree structure and cannot be directly used as the input of Transformer models, GraphCodeBERT allows the model to learn the AST data flow at the cost of losing a considerable part of information. In order to make more effective and comprehensive use of the code structure information provided by AST, UniXcoder [105] improves on the basis of GraphCodeBERT. They construct a one-to-one mapping function to convert an AST into a sequence structure, and then encode it together with the code text. In addition, some models use other ways rather than ASTs to better understand and adapt to the structural characteristics of a code snippet, and perform well. PLBART [106] adds additional denoising strategies to the BART model to correlate PL and NL text. It generates grammatically and logically valid code via large-scale sequence-to-sequence pre-training that incorporates the denoising strategies to better understand program grammars and logical flows. CodeT5 [107] uses the T5 model architecture and makes use of the identifier tagging information and comment information to better model the structural characteristics of code. CodePAD [108] designs an approach based on pushdown automaton (PDA), which takes into account the grammar constraints of programming languages. It ensures the availability of generated code and the correctness of grammar, and avoids the time-consuming and labor-intensive disadvantages of AST-based approaches.
4 Demonstration Systems
To facilitate applications on education scenarios, in this section, we present some off-the-shelf demo cases with brief illustrations of their usages for each task mentioned in Section 3. Information on these demos will be presented in Table V.
For QA, OpenVINO [^10] https://docs.openvino.ai/2023.3/notebooks_section_2_model_demos.html is an open-source toolkit designed for diverse automatic QA, including table, visual, interactive, and LLMs-based QA. PrimeQA [196] is an open-source public repository designed to democratize QA research and simplify the replication of state-of-the-art methods, supporting the training of multilingual QA models with core functionalities like information retrieval, reading comprehension, and question generation.
For MWP, MWPToolkit [197] offers a holistic and extensible framework that provides deep learning-based solvers, popular datasets, and a modular architecture that supports quick replication and innovation of MWP methods. IntMath [^11] https://www.intmath.com/ is an online math platform that combines a mathematical computation engine with artificial intelligence to parse and generate step-by-step natural language answers.
For QC, AutoQG [198] is an online service that provides multilingual question and answer generation, featuring a comprehensive toolkit for model fine-tuning, generation, and evaluation. Besides, Questimator [199] is a system that offers MCQs generation for educational assessment with a certain topic.
For AES and ACS, LinggleWrite [200] is a writing coach tool that offers writing suggestions, proficiency assessments, grammatical error detection and corrective feedback. IntelliMetric [^12] https://www.intellimetric.com/direct/ is an online multilingual AES platform that supports writing evaluation by delivering instant scores with the accuracy and consistency of human experts. Qualified [^13] https://www.qualified.io/ and CoderPad [^14] https://coderpad.io/ are online platforms designed for coding assessments, offering interactive environments to evaluate coding skills.
For GEC and CEC, ALLECS [201] and Grammarly [^15] https://www.grammarly.com/ are web applications that provide online GEC service by offering lightweight GEC systems optimized for users with slow internet connections and comprehensive writing assistance, respectively. CodeGeeX [202] is an open-source and multilingual pre-trained model that enables syntax and function-correct code generation and can be integrated as IDE extensions.
Demo Task Form URL
OpenVINO QA toolkit https://docs.openvino.ai/2023.3/notebooks_section_2_model_demos.html
PrimeQA QA toolkit https://github.com/primeqa
MWPToolkit MWP toolkit https://github.com/LYH-YF/MWPToolkit
IntMath MWP platform https://www.intmath.com/
AutoQG QC application https://autoqg.net/
Questimator QC application https://github.com/methodical-software/questimator
LinggleWrite AES application https://f.linggle.com/
IntelliMetric AES platform https://www.intellimetric.com/direct/
Qualified ACS platform https://www.qualified.io/
CoderPad ACS platform https://coderpad.io/
ALLECS GEC application https://sterling8.d2.comp.nus.edu.sg/allecs/
Grammarly GEC application https://www.grammarly.com/
CodeGeeX CEC extension https://codegeex.cn/
TABLE V: Information on demos that can be applied in education domain.
5 Future Trends
According to the taxonomy and advanced techniques of NLP in the education domain, we discuss several promising future directions for different tasks.
More datasets in education domain. We have summarized the commonly used datasets for educational NLP, which range from question answering to error correction, we also observe there is still a demand for datasets fitting in more diverse education scenarios. Regarding QG tasks, instead of simply considering the single-turn question generation based on the context and answer, we may assume that multi-turn question generation will frequently happen in a tutoring scenario. Specifically, a leading question around a certain knowledge topic is important in guiding students to learn an ultimate subject goal. The distinction between such leading QG and traditional QG challenges lies in the connection between the leading generated questions and the final goal [11]. However, we find there is no existing challenge that has been defined in this way. Regarding AES and GEC tasks, event though we have collected a set of AES and GEC datasets for supporting the researches of natural language learning, there is a lack of AES and GEC datasets in non-English languages, which is fundamental in developing systems for low-resource languages. Furthermore, more interpretable datasets should be proposed to explicit more on the instructional feedback. For example, diverse dimensional explanation could be given in AES datasets and richer error types aligned to the specific grammatical knowledge schema could be provided in GEC datasets. Regarding ACS tasks, public ACS datasets are in high demand as the testbed for evaluating the ACS systems.
The controllable usage of LLMs. Due to the advancement of LLMs, recent studies have started to incorporate LLMs with various educational NLP tasks. Researchers have explored to prompt LLMs for solving zero/few-shot QA, generating questions over knowledge bases [203], and correcting grammatical errors [204] but there remain some drawbacks to be resolved. For example, in QA tasks, even though an LLM is able to act as a “decision center” to understand the questions as well as context and conduct parsing and reasoning, it may perform poorly in a specific knowledge domain. In this case, techniques like Retrieval-Augmented-Generation (RAG) and Multi-Agents are leveraged to enhance the performance of the LLM [33, 48]. But here comes some new questions. For example, the presence of noise or contradictory information retrieving from textbooks can detrimentally affect the quality of LLMs’ outputs. In QG tasks, strongly aligning the outputs of LLMs to the given textbook contexts and generating more controllable questions is an open question. Similarly, in GEC tasks, multiple studies [184] point out that applying LLMs on GEC tasks has a high risk of over-correction, that is failing to preserve the semantics of the original sentences. Therefore, we expect more studies could be proposed to investigate the controllable leverage of LLMs on GEC tasks.
Intervention of difficulty-level control. Most existing NLP techniques for education applications neglect difficulty-level control, which is an important factor in the procedure of learning, comprehension, and assessment at different levels. For example, in QG tasks, generating questions that could be aligned to the level of students is significant in teaching. Even though there are multiple studies [57, 58, 59] trying to model the difficulty level in the QG task, they simply define with the semantic or reasoning complexity of the questions and fail to align the level to the difficulty level of syllabus. For generating distractors, we rarely observe existing studies considering the difficulty level. But this is critical to identify a false candidate answer that is not only relevant to the question but also in accord with the difficulty level of students. In EC tasks, researchers focus more on the accuracy of the correction but neglect the level of the corrected sentences, including the level of substitute words, the syntactic and semantic level of the corrected sentences. Further researches should fuse the difficulty level with various tasks to make it serve educational NLP applications better.
Interpretable educational NLP. Existing studies are able to obtain good results on various tasks like question answering, question construction, automated assessment, and error correction. However, there are limited techniques developed with the perspective of interpretability. As shown in the technical review of QA tasks, recent studies try to explicit the process of thinking in solving Textbook QA and MWPs tasks [24, 127]. In other tasks, interpretability should also be taken into account. It would be promising to show a process of thinking path to make the construction of the questions more rational. For AS and EC tasks, by showing the fine-grained explanation to the scoring and correction, it would not only improve tutoring experience but also benefit the methods as well as systems to diagnose their functions. Even though, we notice a few studies start to develop interpretable GEC systems [174], interpretablity should be further studied and explored with more attention.
Methods with adaptive learning. Even though NLP applications can be applied to education domain, there are a few studies involving adaptive learning to different tasks. For QG task, we notice some researchers [62, 63] attempt to model the learning trajectory of students by incorporating their historical answers to predict their performance on the new questions, which are deemed as a significant feature to determine the difficulty level of the next question. This benefits to build a more customized education system. Such adaptive learning is in high demand to collaborate with other tasks. For example, in EC tasks, when we correct the sentences and codes, we can consider the level of the students based on their input to decide the way to correct. If it is a simple sentence with a mis-used word, we should not substitute it with an obscure words. Similarly, for a piece of erroneous code, we should not correct its errors using the knowledge beyond the knowledge of the students. In this case, developing methods with adaptive learning for various NLP tasks is important.
Integrated systems for education. Even though an individual task can be well-formulated and solved via specific techniques, there is a demand of integrated systems for education. Baladn et al. [205] tuned open-source LLMs for generating teacher responses in BEA 2023 Shared Task. EduChat [206] is a more general LLM in education domain, which is able to automatically assess essays, provide emotional support, and conduct Socratic teaching with the forms of chatting and question answering. In the future, more capabilities like answering textbook questions, automated quiz making, should be integrated to empower the system, making it perform as an efficient teaching assistant. Meanwhile, an integrated could ensemble the tutoring abilities in the dimension of different subjects such as math, computer science, linguistics and so on. In addition, exploring more interactive ways between the users and machines of an integrated system is worth studying. Generating responses from the system based on the question from the users may be not practical. In real-scenario, automatic answer or instruction prompting instead of waiting for the questions from students would help provide teaching assistance without raising up a question in purpose.
6 Conclusions
This survey attempted to provide a comprehensive overview of NLP in the education domain, as depicted in Figure 2. We highlight that NLP techniques can be applied to various procedure of education. After process and analysis the raw data from diverse data source, NLP applications can be directly serve to tutoring and learning. Our survey highlights four representative tasks (i.e., question answering, question construction, automated assessment, and error correction) with their fine-grained sub-tasks as the focus of this survey due to their significant impact in teaching and learning. According to the taxonomy, we go deep into these tasks and illustrate the evolution of their techniques. For each task, we describe its task definition, unique challenges in education domain, and its corresponding solutions. Early studies developed based on rule-based, deep learning-based methods to solve these tasks. Thank to the advancement of LLMs, a number of methods start collaborating LLMs with traditional methods. This innovation has led to a easier paradigm that improve the performance and decrease the demand of the annotation data. However, LLMs also encounter some new issues like domain knowledge incorporation and over-correction when we apply them to specific tasks. Therefore, future studies need to fuse LLMs to the diverse tasks with the consideration of solving the their drawbacks. Then, to help readers apply applications on education scenarios, we present the off-the-shelf toolkits and demonstrations. Despite the strides made in NLP applications on education domain, there still remains improvement room for diverse tasks. In the last section, we suggest some future directions for NLP in the education domain.
We would like to note that besides the highlighted tasks in this survey, there are some other tasks that can contribute to the educational scenarios. For example, knowledge graph construction [12, 207, 208] aims to organize and extract the knowledge points from the unstructured data sources to form a knowledge graph, which is able to demonstrate the connection between the schema for various subjects and denote to the downstream applications. Moreover, another NLP application controllable text generation [209, 210, 211] aims to generate text with given style strength, length, or components, which can help teachers prepare sentences teaching materials for the linguistics courses. In summary, the demand for educational applications is increasing and educational NLP will thrive to be a promising research area. We hope this survey will give a comprehensive picture of educational NLP and we encourage more contribution in this field.